Air is being pumped into a spherical balloon at a rate of 100cm3/min. Determined the rate at which the radius of the balloon is increasing when the radius of the balloon is 10cm?
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Air is being pumped into a spherical balloon at a rate of 100cm3/min. Determined the rate at which the radius of the balloon is increasing when the radius of the balloon is 10cm?
------------------
V = 4pi*r^3/3
dV/dt = 4pi*r^2*dr/dt = 100
dr/dt = 25/(pi*r^2) = 1/(4pi) cm/minute
===========================
This solution ignores physical properties such as changes in temperature and pressure of the gas.

Answer by ikleyn(52787)

.

Actually, this problem is DEFECTIVE.

The fact that that the air is being pumped into a spherical balloon at a rate of 100cm3/min,
physically DOES NOT MEAN that the rate of increasing the balloon volume is the same.


Thinking opposite means making a rude physical mistake.


Lowest possible score to the problem's composer for his (or her) unsatisfactory work.